Charles has decided to open a​ lawn-mowing company. To do​ so, he purchases mowing equipment for ​$6,000​, buys gasoline ​($1.50 in gas is required to mow each​ yard), and pays a helper ​$20.00 per yard. Prior to opening the lawn​ company, Charles earned ​$5,000 as a lifeguard at the neighborhood swimming pool. Assume the money he used to purchase the mowing equipment could otherwise have earned 1 percent per year in the bank and that the mowing equipment depreciates at 20 percent per year. Charles plans to mow 250 yards per year. What is​ Charles's implicit cost LOADING... of​ production? ​Charles's implicit cost of production is ​$ nothing per year. ​(Enter your response as an​ integer.)
Answer #1

There is basically two types of cost, implicit cost, and explicit cost.

Implicit cost is the opportunity cost or the foregone value of an alternative.

Explicit cost is the actual cost or the direct payment in the process of production. The cost of equipment, helper and gasoline is an explicit cost. Depreciation is also an explicit cost.

So here implicit cost is the sum of earning foregone as a lifeguard and foregone interest rate.

Implicit cost = $5000 + 1% of $6,000 = $5060.
Answer #2

Implicit cost is the cost incurred on the part of the owner for forgoing the next best alternatives. or the cost of self-employed resources and self-owned resources. 

Here, Charles has forgone the interest on his capital, that he could have otherwise earned, and his income of lifeguard job given up to open a lawn mowing company. 


Therefore, the implicit cost = 5000 +6% of 6000 = 5000+60 = 5060

source: Books
answered by: Sana Rajar
Answer #3

The implicit cost will be the amount from lifeguarding + the amount from the interest + the amount after the first depreciation.


Implicit cost = $5000 + 1% of $6000 + 20% of $6000

= 5000 + 60 + 1200

= $6260

answered by: alexawxyz